OpenStudy (anonymous):

Solve |2x + 8| = 6. Show its equivalent compound sentence and show its solution. One point for each.

OpenStudy (anonymous):

\[2x+8=6\] or \[2x+8=-6\] solve for \(x\)

OpenStudy (anonymous):

use 2x + 8= 6 and 2x + 8=-6 x=1 or -7

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

\(x=-7\) looks good, but not \(x=1\)

OpenStudy (anonymous):

\(2x+8=6\iff 2x=-2\iff x=-1\)